# RLVR-World: Training World Models with Reinforcement Learning

### v1

@@ 5.2 Web Page State Prediction
We further evaluate our approach on more realistic web navigation scenarios, using a web page state transition dataset collected by WMA [8] from the WebArena benchmark [69]. For training and testing, we select a 7K-sample subset consisting of shorter-length samples to avoid out-of-memory issues during training. In this task, a website’s state is represented by its accessibility tree, which is simplified from its Document Object Model (DOM) tree, and an LLM is used to predict state transitions after user actions such as clicking. Item changes in the accessibility tree caused by actions are extracted using the Hungarian algorithm, enabling the model to predict these changes directly. Unlike WMA, which generates natural language descriptions of state changes, this design choice facilitates clear verification during the RLVR stage.
@@
Following the previous setup, we adopt DeepSeek-R1-Distill-Qwen-1.5B as our base model. We first perform supervised fine-tuning (SFT) using chain-of-thought (CoT) data provided by WMA [8]. Subsequently, we apply RLVR, using the F1 score between predicted item changes $\Delta\hat{s}$ and ground truth $\Delta s$ as the reward function: $R=\operatorname{F1}(\Delta\hat{s},\Delta s)$. The F1 score, defined as the harmonic mean of precision and recall, is computed by treating precision as the proportion of correctly predicted item changes among all generated ones, and recall as the proportion of correct predictions relative to the ground-truth item changes. An item change is considered correct only if it exactly matches the corresponding ground truth.
@@ 5.3 Application: Model Predictive Control for Web Agents
Following WMA [8], we build web agents for the WebArena benchmark [69], composed of three components: a policy model, a world model, and a value model. The model predictive control pipeline proceeds as follows: the policy model first proposes multiple candidate actions; the world model then predicts the outcomes of these actions; finally, the value model scores each predicted outcome based on the task goal. The action with the highest score is selected for execution. For both the policy and value models, we use DeepSeek-V3 [33], while the world model is taken from our trained models in the previous section. Additional details are provided in Appendix A.3.
@@ Results.
6 Evaluating Video World Models with RLVR
6.1 Setup
Dataset and task.
We use the RT-1 robotic manipulation dataset [5], which contains 87,212 tabletop teleoperation trajectories collected from a Google Robot, with 1% left for testing. Each frame of visual observation has a resolution of $256\times 320$, and the action space consists of 13 dimensions, including arm and base movement. We evaluate world models on two task settings: (1) Single-step prediction, formulated as $p(s_{t+1}\mid s_{t-3:t},a_{t-3:t})$: predicting the next observation given the past four-step observations and actions; (2) Multi-step prediction, formulated as $p(s_{t+1:t+7}\mid s_{t},a_{t:t+6})=\prod_{i=t+1}^{t+7}p(s_{i}\mid s_{t:i-1},a_{t:i-1})$: predicting next seven-step observation conditioed on the current observation and future action sequence. Predictions are assessed against ground-truth observations using frame-level metrics: MSE, PSNR [24], SSIM [58], and LPIPS [68].
World model.
Since no off-the-shelf video world models are available, we pre-train variants of iVideoGPT [61] as base models by ourselves. For each trajectory segment, observations and actions are tokenized and concatenated into a unified token sequence. We train an image tokenizer to independently tokenize video frames for single-step prediction, but train a compressive tokenizer from iVideoGPT to mitigate sequence length explosion for multi-step prediction. Each action dimension is discretized into 256 uniform bins, determined based on its value range across the entire dataset. During RLVR fine-tuning, we define the reward function as the sum of L1 and perceptual loss between decoded predicted and ground-truth frames: $R=-\sum_{\tau=t+1}^{t+T}\left[L_{1}(\hat{s}_{\tau},s_{\tau})+\operatorname{LPIPS}(\hat{s}_{\tau},s_{\tau})\right]$, commonly used in visual tokenizer training [10]. See implementation details in Appendix A.4.
Table 3: Video world model: robot manipulation trajectory prediction. We report the mean and standard deviation for each metric calculated over three sampling runs. MSE, LPIPS, and SSIM scores are scaled by 100 for better readability. $\Delta$: relative performance gains from RLVR.
@@ 6.1 Setup
Figure 3: Learning curves of video world models. Note the significant difference in the $x$-axis scale between the pre-training and post-training stages.
@@ 6.2 Main Results
As shown in Table 3, RLVR-World significantly improves the base model across all visual metrics, demonstrating more accurate and perceptually better video predictions, showcased in Figure 6. We highlight that these gains are achieved with only hundreds of RLVR gradient steps, compared to hundreds of thousands required for MLE pre-training (see the training curves in Figure 3). Even after continuing pre-training for multi-step prediction with 150k additional steps–1000$\times$ more than RLVR fine-tuning–the resulting LPIPS score remains at 14.5, still substantially lagging behind.
@@ 6.3 Model Analysis
Prior studies [31] have identified the likelihood objective as a primary cause of repetition in LLM generation. We observe a similar phenomenon in multi-step video prediction, as showcased in Figure 6. This likely stems from the fact that approximately 20% of tokens in each frame remain unchanged from the previous frame, encouraging the model to exploit this shortcut. By directly optimizing video-level prediction metrics rather than next-token likelihood, RLVR effectively mitigates this issue, reducing the repetition rate from 48.6% to 9.9%. To ensure our improvements are not merely due to reducing repetition, in Table 3, we include an additional baseline that repeatedly queries the base model for predictions until a non-repetitive output is sampled.
@@
To further demonstrate the effect of direct metric optimization, we post-train five variants of the base model using five different metrics as reward functions: MAE, MSE, PSNR, SSIM, and LPIPS. As shown in Figure 4, models fine-tuned with a specific metric generally achieve the best performance when evaluated on that same metric.
@@
We evaluate the test-time scaling behavior of our base and RLVR-trained models by reporting the best metric achieved across $N$ samples [36] in Figure 4. RLVR-World improves one-shot performance, even outperforming the base model’s best-of-$5$ results. This is particularly valuable in practical scenarios where generating large numbers of samples is computationally expensive and ground-truth comparisons are unavailable. However, as $N$ increases to 100, the base model catches up and eventually surpasses RLVR-trained models, echoing findings from Yue et al. [67]. This suggests limitations of current RLVR methods and ample opportunities for future research.
@@
Figure 4: Model analysis. (a) Test-time scaling: best performance among different numbers of generated samples. (b) RL training scaling: learning curves using different group sizes in GRPO. (c) Metric-oriented optimization: RLVR-World trained and tested on different visual metrics.
@@ 6.4 Application: Real2Sim Policy Evaluation
Following SIMPLER [32], we evaluate four policy checkpoints from RT-1 [5] and RT-1-X [41] on six tasks involving opening and closing top, middle, and bottom drawers. Starting from a real-world observation frame, policies can interact with video world models to roll out neural simulated trajectories, allowing for policy evaluation without real-world deployment. Since our preliminary attempts on VLM-based automatic evaluation [53] fail to provide reliable judgments, we rely on human annotators to assess the success of simulated trajectories. Besides our base model, we compare against the simulators developed in SIMPLER. Experimental details can be found in Appendix A.5.
@@ 7 Discussion and Limitations
In this work, we pioneer RLVR for training world models across language and video modalities, with practical applications in web navigation and robotic manipulation. We believe RLVR has the potential to become a general post-training paradigm for generative models. To this end, several challenges remain for future exploration: (1) Task-aligned rewards: While classical visual metrics align better with the world modeling task than MLE, they still fail to fully capture user-intended qualities. Incorporating constraints such as physical rules and temporal consistency will require more sophisticated reward designs. (2) Breaking performance barriers: Although RLVR yields significant gains, training typically converges within hundreds of steps. Unlocking continual improvements calls for deeper analysis of bottlenecks in models, data, and algorithms; (3) Out-of-distribution (OOD) generalization: Inspired by RLVR’s success in enabling LLMs to generalize beyond training domains [49], it is important to study whether similar benefits extend to world models, particularly for counterfactual reasoning on OOD actions in sequential decision-making.
@@ A.1 Text Game State Prediction
ByteSized32-State-Prediction [57] comprises 76,369 transitions collected from 31 distinct text-based games. A curated subset of 2,954 high-quality transitions forms the official test set, which we use for both our SFT and RLVR experiments.
@@
Each state is represented as a JSON object, describing the environment objects and their associated properties (e.g., a dish with a "clean" or "dirty" status). Actions are described in natural language (e.g., “clean the dish”), and the world simulator aims to predict the resulting state, the task reward based on predefined rules (e.g., $+1$ for each clean dish), and task completion (e.g., when all dishes are clean). We categorize samples into "unchanged cases" when the action leaves the state unchanged, i.e., the ground truth next state is exactly the same as the current state, and "changed cases" otherwise.
To reduce output complexity, as suggested by the original paper [57], we train our world models for state-difference prediction, generating updates solely for objects whose properties have changed. This design mitigates challenges in generating long, correctly formatted outputs with our 1.5B small base model. Prompt examples are provided in Appendix D.
@@
To overcome the poor performance of our base model, DeepSeek-R1-Distill-Qwen-1.5B [^3] DeepSeek-R1-Distill-Qwen-1.5B follows the MIT License., especially its 0.08% accuracy on changed cases (see Table 1), we first apply supervised fine-tuning (SFT) with LoRA [23] to enable effective RL training.
@@
SFT Batch size $16$ $8$
LoRA rank [23] $32$ $32$
@@
Learning rate $1\times 10^{-5}$ $1\times 10^{-5}$
@@
KL loss coefficient $1\times 10^{-3}$ $1\times 10^{-3}$
Group size $5$ $5$
Learning rate $1\times 10^{-6}$ $1\times 10^{-6}$
Sampling Top-$p$ $1.0$ $1.0$ for prediction $0.95$ for MPC
@@ A.2 Web Page State Prediction
For our new task formulation, we construct target responses for each sample in the WMA dataset by concatenating the chain-of-thought (CoT) annotations, which are generated by GPT-4o-mini and released by the authors, and our extracted next-state changes. We then perform supervised fine-tuning (SFT) with LoRA [23] on our base model, DeepSeek-R1-Distill-Qwen-1.5B. Training hyperparameters are summarized in Table 4.
@@
Precision ($\text{Precision }(\mathrm{P})$ =$=\begin{cases}\frac{\text{TP}}{|\Delta\hat{s}|},&\text{if }|\Delta\hat{s}|>0\\ 1,&\text{if }|\Delta\hat{s}|=0\text{ and }|\Delta s|=0\\ 0,&\text{if }|\Delta\hat{s}|=0\text{ and }|\Delta s|>0\end{cases}$
Recall ($\text{Recall }(\mathrm{R})$ =$=\begin{cases}\frac{\text{TP}}{|\Delta s|},&\text{if }|\Delta s|>0\\ 1,&\text{if }|\Delta\hat{s}|=0\text{ and }|\Delta s|=0\\ 0,&\text{if }|\Delta\hat{s}|>0\text{ and }|\Delta s|=0\end{cases}$
[…]

### v2

@@ 5.2 Web Page State Prediction
We further evaluate our approach on more realistic web navigation scenarios, using a web page state transition dataset collected by WMA chae2024web from the WebArena benchmark zhou2023webarena . For training and testing, we select a 7K-sample subset consisting of shorter-length samples to avoid out-of-memory issues during training. In this task, a website’s state is represented by its accessibility tree, which is simplified from its Document Object Model (DOM) tree, and an LLM is used to predict state transitions after user actions such as clicking. Item changes in the accessibility tree caused by actions are extracted using the Hungarian algorithm, enabling the model to predict these changes directly. Unlike WMA, which generates natural language descriptions of state changes, this design choice facilitates clear verification during the RLVR stage.
@@
Following the previous setup, we adopt DeepSeek-R1-Distill-Qwen-1.5B as our base model. We first perform supervised fine-tuning (SFT) using chain-of-thought (CoT) data provided by WMA chae2024web . Subsequently, we apply RLVR, using the F1 score between predicted item changes $\Delta\hat{s}$ and ground truth $\Delta s$ as the reward function: $R=\operatorname{F1}(\Delta\hat{s},\Delta s)$. The F1 score, defined as the harmonic mean of precision and recall, is computed by treating precision as the proportion of correctly predicted item changes among all generated ones, and recall as the proportion of correct predictions relative to the ground-truth item changes. An item change is considered correct only if it exactly matches the corresponding ground truth.
@@ 5.3 Application: Model Predictive Control for Web Agents
Following WMA chae2024web , we build web agents for the WebArena benchmark zhou2023webarena , composed of three components: a policy model, a world model, and a value model. The model predictive control pipeline proceeds as follows: the policy model first proposes multiple candidate actions; the world model then predicts the outcomes of these actions; finally, the value model scores each predicted outcome based on the task goal. The action with the highest score is selected for execution. For both the policy and value models, we use DeepSeek-V3 liu2024deepseek , while the world model is taken from our trained models in the previous section. Additional details are provided in Appendix A.3.
@@ Results.
Table 3: Video world model: robot manipulation trajectory prediction on RT-1. We report the mean and standard deviation for each metric calculated over three sampling runs. MSE, LPIPS, and SSIM scores are scaled by 100 for better readability. $\Delta$: relative performance gains from RLVR.
@@
RLVR-World
(w/ rep. penalty reward)
0.0% 0.506
$\pm$0.002
24.0
$\pm$0.02
82.2
$\pm$0.01
13.7
$\pm$0.02
Figure 3: Learning curves of video world models on RT-1. Note the significant difference in the $x$-axis scale between the pre-training and post-training stages.
6 Evaluating Video World Models with RLVR
6.1 Setup
Dataset and task.
We primarily use the RT-1 robotic manipulation dataset brohan2022rt for our experiments, which contains 87,212 tabletop teleoperation trajectories collected from a Google Robot, with 1% left for testing. Each frame of visual observation has a resolution of $256\times 320$, and the action space consists of 13 dimensions, including arm and base movement. To compare with state-of-the-art models, we also include tabletop pushing (PushT) chi2023diffusion and deformable object manipulation (Rope and Granular) zhang2024adaptigraph datasets from DINO-WM zhou2024dino . We evaluate world models on two task settings: (1) Single-step prediction, formulated as $p(s_{t+1}\mid s_{t-T+1:t},a_{t-T+1:t})$: predicting the next observation given the past $T$-step observations and actions; (2) Multi-step prediction, formulated as $p(s_{t+1:t+T}\mid s_{t},a_{t:t+T-1})=\prod_{i=t+1}^{t+T}p(s_{i}\mid s_{t:i-1},a_{t:i-1})$: predicting next $T$-step observation conditioed on the current observation and future action sequence. Predictions are assessed against ground-truth observations using frame-level metrics: MSE, PSNR huynh2008scope , SSIM wang2004image , and LPIPS zhang2018unreasonable .
World model.
Since no off-the-shelf general-purpose video world models are available, we pre-train variants of iVideoGPT wu2024ivideogpt on target datasets as base models by ourselves. For each trajectory segment, observations and actions are tokenized and concatenated into a unified token sequence. We train an image tokenizer to independently tokenize video frames for single-step prediction, but train a compressive tokenizer from iVideoGPT to mitigate sequence length explosion for multi-step prediction. Each action dimension is discretized into 256 uniform bins, determined based on its value range across the entire dataset. During RLVR fine-tuning, we define the reward function as the sum of L1 and perceptual loss between decoded predicted and ground-truth frames: $R=-\sum_{\tau=t+1}^{t+T}\left[L_{1}(\hat{s}_{\tau},s_{\tau})+\operatorname{LPIPS}(\hat{s}_{\tau},s_{\tau})\right]$, commonly used in visual tokenizer training esser2021taming . See implementation details in Appendix A.4.
Table 4: Video world model: comparisons with state-of-the-art on PushT, Rope, and Granular. We follow the single-step prediction setting and include the baseline results reported in DINO-WM zhou2024dino . In addition, we report our reproduced evaluation results using the public DINO-WM checkpoint on PushT. LPIPS and SSIM scores are scaled by 100.
PushT Rope Granular
[…]
Recurrent latent space models
R3M nair2022r3m ; zhou2024dino 4.5 95.6 2.3 98.2 8.0 91.7
ResNet he2016deep ; zhou2024dino 6.3 95.0 2.5 98.0 8.0 91.5
DINO CLS oquab2023dinov2 ; zhou2024dino 3.9 97.3 2.9 98.0 8.6 91.2
DINO-WM
zhou2024dino
0.7 98.5 0.9 98.5 3.5 94.0
DINO-WM
(Public checkpoint)
zhou2024dino
3.39 96.38 - - - -
Diffusion models
AVDC ko2023learning 4.6 95.9 6.0 97.9 10.6 90.9
Autoregressive models
Base
0.83 98.28 3.03 97.86 3.14 94.79
RLVR-World
0.70 98.46 2.08 98.14 2.42 95.42
@@ 6.2 Main Results
[…]
Comparison with state-of-the-art.
We then compare our models against advanced world models, including all recurrent latent state models introduced by DINO-WM zhou2024dino , which use different latent spaces from pre-trained encoders such as R3M, ResNet, and DINOv2, and an action-conditioned diffusion model, AVDC ko2023learning . As shown in Table 4, after RLVR, our model achieves overall performance comparable to the strongest baseline, DINO-WM, and surpasses it by a significant margin on the most challenging particle-based dataset, Granular. Additional model predictive control results on PushT are provided in Appendix A.4.2.
@@ 6.3 Model Analysis
Prior studies li2023repetition have identified the likelihood objective as a primary cause of repetition in LLM generation. We observe a similar phenomenon in multi-step video prediction, as showcased in Figure 6. This likely stems from the fact that approximately 20% of tokens in each frame remain unchanged from the previous frame, encouraging the model to exploit this shortcut. By directly optimizing video-level prediction metrics rather than next-token likelihood, RLVR effectively mitigates this issue, reducing the repetition rate from 48.6% to 9.9%. To ensure our improvements are not merely due to reducing repetition, in Table 3, we include an additional baseline that repeatedly queries the base model for predictions until a non-repetitive output is sampled.
@@
To further demonstrate the effect of direct metric optimization, we post-train five variants of the base model using five different metrics as reward functions: MAE, MSE, PSNR, SSIM, and LPIPS. As shown in Figure 4, models fine-tuned with a specific metric generally achieve the best performance when evaluated on that same metric. Additionally, to show the effectiveness of our method for non-differentiable rewards and achieve zero repetition rate, we introduce an additional repetition penalty reward, defined as the negative rate of consecutive identical frames. In Table 3, after training the model with this extended reward, we observe that it maintains comparable prediction performance while effectively eliminating repetition artifacts.
@@
We evaluate the test-time scaling behavior of our base and RLVR-trained models by reporting the best metric achieved across $N$ samples ma2025inference in Figure 4. RLVR-World improves one-shot performance, even outperforming the base model’s best-of-$5$ results. This is particularly valuable in practical scenarios where generating large numbers of samples is computationally expensive and ground-truth comparisons are unavailable. However, as $N$ increases to 100, the base model catches up and eventually surpasses RLVR-trained models, echoing findings from Yue et al. yue2025does . This suggests limitations of current RLVR methods and ample opportunities for future research.
@@
Figure 4: Model analysis on RT-1. (a) Test-time scaling: best performance among different numbers of generated samples. (b) RL training scaling: learning curves using different group sizes in GRPO. (c) Metric-oriented optimization: RLVR-World trained and tested on different visual metrics.
@@ 6.4 Application: Real2Sim Policy Evaluation
Following SIMPLER li2024evaluating , we evaluate four policy checkpoints from RT-1 brohan2022rt and RT-1-X padalkar2023open on six tasks involving opening and closing top, middle, and bottom drawers. Starting from a real-world observation frame, policies can interact with video world models to roll out neural simulated trajectories, allowing for policy evaluation without real-world deployment. Since our preliminary attempts on VLM-based automatic evaluation team2024gemini fail to provide reliable judgments, we rely on human annotators to assess the success of simulated trajectories. Besides our base model, we compare against the simulators developed in SIMPLER. Experimental details can be found in Appendix A.5.
@@ 7 Discussion and Limitations
In this work, we pioneer RLVR for training world models across language and video modalities, with practical applications in web navigation and robotic manipulation. We believe RLVR has the potential to become a general post-training paradigm for generative models. To this end, several challenges remain for future exploration. From the algorithmic perspective: (1) Breaking performance barriers: Although RLVR yields significant gains, training typically converges within hundreds of steps. Unlocking continual improvements calls for deeper analysis of bottlenecks in models, data, and algorithms; (2) Out-of-distribution (OOD) generalization: Inspired by RLVR’s success in enabling LLMs to generalize beyond training domains shen2025satori , it is important to study whether similar benefits extend to world models, particularly for counterfactual reasoning on OOD actions in sequential decision-making. For developing foundation world models: (1) Post-training general-purpose world models: Our current video world models adopt a two-stage training process on the same dataset. We believe that the full potential of RLVR will be further unlocked once the community develops general-purpose video world models agarwal2025cosmos ; li2025unified . This would enable the complete paradigm of supervised pre-training on diverse domains $\rightarrow$ supervised fine-tuning $\rightarrow$ reinforcement fine-tuning, which has already proven effective in our language world model experiments. (2) Broader application across model classes: While the core insight behind RLVR-World is model-agnostic, GRPO algorithms for diffusion models xue2025dancegrpo ; liu2025flow are concurrently developed and thus fall outside the scope of our current study. (3) Task-aligned rewards: While classical visual metrics align better with the world modeling task than MLE, they still fail to fully capture user-intended qualities. Incorporating constraints such as physical rules and temporal consistency will require more sophisticated reward designs.
Acknowledgements
We would like to thank many colleagues, in particular Chaoyi Deng and Haixu Wu, for their valuable discussion. This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
@@ A.1 Text Game State Prediction
ByteSized32-State-Prediction wang2024can comprises 76,369 transitions collected from 31 distinct text-based games. A curated subset of 2,954 high-quality transitions forms the official test set, which we use for both our SFT and RLVR experiments.
@@
Each state is represented as a JSON object, describing the environment objects and their associated properties (e.g., a dish with a "clean" or "dirty" status). Actions are described in natural language (e.g., “clean the dish”), and the world simulator aims to predict the resulting state, the task reward based on predefined rules (e.g., +1+1 for each clean dish), and task completion (e.g., when all dishes are clean). We categorize samples into "unchanged cases" when the action leaves the state unchanged, i.e., the ground truth next state is exactly the same as the current state, and "changed cases" otherwise.
To reduce output complexity, as suggested by the original paper wang2024can , we train our world models for state-difference prediction, generating updates solely for objects whose properties have changed. This design mitigates challenges in generating long, correctly formatted outputs with our 1.5B small base model. Prompt examples are provided in Appendix D.
@@
To overcome the poor performance of our base model, DeepSeek-R1-Distill-Qwen-1.5B [^3] DeepSeek-R1-Distill-Qwen-1.5B follows the MIT License., especially its 0.08% accuracy on changed cases (see Table 1), we first apply supervised fine-tuning (SFT) with LoRA hu2022lora to enable effective RL training.
@@
SFT Batch size $16$ 88
LoRA rank hu2022lora $32$ $32$
@@
Learning rate $\times 10^{-5} 1×1$$\times 10^{-5}$
@@
KL loss coefficient $\times 10^{-3} 1×1$$\times 10^{-3}$
Group size 55 55
Learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$
Sampling Top-pp $1.0$
$1.0$ for prediction
$0.95$ for MPC
@@ A.2 Web Page State Prediction
For our new task formulation, we construct target responses for each sample in the WMA dataset by concatenating the chain-of-thought (CoT) annotations, which are generated by GPT-4o-mini and released by the authors, and our extracted next-state changes. We then perform supervised fine-tuning (SFT) with LoRA hu2022lora on our base model, DeepSeek-R1-Distill-Qwen-1.5B. Training hyperparameters are summarized in Table 5.
@@
$\text{Precision }(\mathrm{P})$ $=\begin{cases}\frac{\text{TP}}{|\Delta\hat{s}|},&\text{if }|\Delta\hat{s}|>0\\ 1,&\text{if }|\Delta\hat{s}|=0\text{ and }|\Delta s|=0\\ 0,&\text{if }|\Delta\hat{s}|=0\text{ and }|\Delta s|>0\end{cases}$
$\text{Recall }(\mathrm{R})$ $=\begin{cases}\frac{\text{TP}}{|\Delta s|},&\text{if }|\Delta s|>0\\ 1,&\text{if }|\Delta\hat{s}|=0\text{ and }|\Delta s|=0\\ 0,&\text{if }|\Delta\hat{s}|>0\text{ and }|\Delta s|=0\end{cases}$
$\text{F1 Score (Reward }R\text{)}$ $=\begin{cases}\frac{2\cdot\mathrm{P}\cdot\mathrm{R}}{\mathrm{P}+\mathrm{R}},&\text{if }\mathrm{P}+\mathrm{R}>0\\ 0,&\text{otherwise}\end{cases}$